OpenStudy (anonymous):

how to re arrange the following as y=mx+c 1. 1/x+5=2/y 2. 3/x=1/y

OpenStudy (whpalmer4):

\[\frac1{x + 5} = \frac2y\] Is that correct? You really should either use the Equation editor or parentheses to make the meaning of your equations clear!

OpenStudy (campbell_st):

just cross multiply so the 1st equation becomes \[y = 2(x + 5)\] now just expand and simplify to the equation to be in the form of y = mx + c same process for the 2nd equation.

OpenStudy (anonymous):

1st question i understood but second one i just have no idea to solve it as y=mx+b

OpenStudy (whpalmer4):

\[\frac3x=\frac1y\]Multiply both sides by \(xy\) and you get \[3y = x\]\[y=\frac x3\]

OpenStudy (whpalmer4):

which is \[y=(\frac{1}3)x+b\]where \(b=0\)

OpenStudy (anonymous):

thanq dear